# Training-Free Safe Denoisers for Safe Use of Diffusion Models

### v2

@@ B.2 Additional Results
We present additional qualitative results across three experimental scenarios: (1) Text-to-Image Generation for preventing nudity, (2) Sexual Debiasing in unconditional generation for facial images, and (3) Class-Conditional Generation, where reference images serve as constraints not to generate. To systematically demonstrate the effectiveness of our approach, we present the results in sequence, beginning with text-to-image generation followed by unconditional generation and concluding with conditional generation.
Figure B.3: Generated images by baselines and ours on Ring-A-Bell (Tsai et al., 2024)
Figure B.4: Generated images by baselines and ours on UnlearnDiff (Zhang et al., 2024b)
Figure B.5: Uncurated generated images by SAFREE+Ours on CoCo30K
@@
Figure B.6: Comparison of Safe Denoiser against existing approaches when negation on female.
Figure B.7: Comparison of Safe Denoiser against existing approaches when negation on Chihuahua. This comparison includes non-dog related ImageNet classes, which include Tench, Garbage Truck, Church, Spoonbill, and Great White Shark.
Figure B.8: Classifier guidance diffusion model generated samples when negating on Chihuahua. This comparison includes non-dog-related ImageNet classes mentioned in B.7 along with the dog-related classes in Figure  5 which are Pomeranian, Yorkshire Terrier, and Shih Tzu.
Figure B.9: Sparse Repellency generated samples when negating on Chihuahua. The same classes are selected as  B.8.
Figure B.10: Safe Denoiser generated samples when negating on Chihuahua. The same classes are selected as  B.8.

### v4

@@ E.2 Qualitative Results
We present additional qualitative results across three experimental scenarios: (1) Text-to-Image Generation for preventing nudity and inappropriate images, (2) Sexual Debiasing in unconditional generation for facial images, and (3) Class-Conditional Generation, where reference images serve as constraints not to generate. To systematically demonstrate the effectiveness of our approach, we present the results in sequence, beginning with text-to-image generation followed by unconditional generation and concluding with conditional generation. To facilitate straightforward understanding, we include as many figures and qualitative comparisons as possible.
E.3 Text-to-Image Generation
Safe Generation against Nudity Prompts
We present a qualitative comparison across baselines and ours. All figures are generated using the same text prompts. We decide to exclude the case of MMA-Diffusion since prompts in this dataset generate pornographic images by baselines, which is not suitable for academic purpose. Hence, we select text prompts from Ring-A-Bell [37] and UnlearnDiff [36]. From Figure E.4 to Figure E.5, we observe that our model effectively eliminates nudity information while preserving textual information.
Figure E.4: Generated images by baselines and ours on Ring-A-Bell [37]
Figure E.5: Generated images by baselines and ours on UnlearnDiff [36]
Inappropriate Probability in CoPro Dataset
In our evaluation on the CoPro dataset [42], we apply our method with images from the I2P dataset [2], which includes a wide range of sensitive categories: {hate, harassment, violence, self-harm, sexual content, shocking content, and illegal activities}. Among these, we focus on the ’Self-Harm’ category. Self-harm content is suitable for graphical illustration, distinguishing it an appropriate and interpretable case for homogenous visual inspection in the public domain. Unlike other categories—where perceptions of appropriateness can vary widely across cultural and personal contexts—’Self-harm’ is typically associated with somber or distressing imagery that is broadly and publicly recognized as unsafe.
Figure E.6: Generated images by baselines and ours on CoPro [42]. All texual prompts are labeled as ’Self-Harm’. This dataset also provides safe alternatives, and we present both.
As illustrated in Figure E.6, our Safe Denoiser effectively reduces the generation of implicit unsafe content while preserving coherence with the provided prompts. This underscores its ability to both detect and suppress sensitive content without compromising the semantic alignment of textual prompts. From this figure, it is evident that negative prompts do not yield significant results in mitigating the generation of sad and gloomy atmospheres, particularly for conveying feelings of collapse. Conversely, our Safe Denoiser tends to generate images that more accurately reflect the literal meanings of the textual prompts. This tendency contributes to a reduction in the likelihood of generating content that evokes feelings of ’Self-harm’.
Alignment of Textual Prompts
We present uncurated generated images from the CoCo dataset. This dataset encompasses a wide range of textual prompts that cover various lengthy and diverse situations. As shown in Figure E.7 and Table  2, we conclude that our approach does not compromise the performance of generating normal images. Instead, it focuses on addressing the challenge of generating potentially unsafe images.
Figure E.7: Uncurated generated images by SAFREE+Ours on CoCo30K
E.4 Unconditional Generation: Sexual Debiasing
We present uncurated generated images created by the DM trained on the FFHQ dataset [49]. This dataset lacks explicit labels indicating sexual information. However, we observe a tendency for this model to generate female images more frequently than male images, as shown in Table  6. On the other hand, when we utilize Safe denoiser with female images, we mitigate the potential bias towards female images and achieves generating images uniformly distributed across sexual information. Figure E.8 illustrates that the generated images align with the numerical results presented in Table  6.
@@ E.4 Unconditional Generation: Sexual Debiasing
Figure E.8: Comparison of Safe Denoiser against existing approaches when negation on female.
E.5 Class-conditional Generation : Negation of Specific Class
We present uncurated images that focus on negating a specific Chihuahua class. Here are two experimental setups. First, we employ class conditional guidance on the ‘Chihuahua’ class and simultaneously use the Safe denoiser to work with negative images sampled from the ‘Chihuahua’ class in the validation split. We observe that the SR does not follow homogeneous images that align with the superclass, ‘Dog’, but our method produces similar small dogs but not matched with ‘Chihuahua’ as shown in Figure E.9.
Second, we qualitatively evaluate that our method with negative images from the ‘Chihuahua’ class works when class guidance is applied to classes other than ‘Dogs’, for example, ‘Tench’ and ‘Truck’. This result is shown in Figure E.10. We observe that the SR sometimes depicts different classes even when class guidance is applied, but our method aligns with homogeneous classes following class guidance even when the Safe denoiser works with ‘Chihuahua’ images. This indicates that our method effectively tackles specific concepts and preserves the original performance when it is not mutually correlated.
Figure E.9: Generated samples when negating the Chihuahua class, primarily producing visually similar small dog breeds.
Figure E.10: Comparison of Safe Denoiser against existing approaches when negation on Chihuahua. This comparison includes non-dog related ImageNet classes, which include Tench, Garbage Truck, Church, Spoonbill, and Great White Shark.
Additional graphical illustrations are presented in the following figures from Figure E.11 to Figure E.13.
Figure E.11: Classifier guidance diffusion model generated samples when negating on Chihuahua. This comparison includes non-dog-related ImageNet classes mentioned in E.10 along with the dog-related classes in Figure  E.9 which are Pomeranian, Yorkshire Terrier, and Shih Tzu.
Figure E.12: Sparse Repellency generated samples when negating on Chihuahua. The same classes are selected as  E.11.
Figure E.13: Safe Denoiser generated samples when negating on Chihuahua. The same classes are selected as  E.11.